OpenStudy (anonymous):

Can some one help me multiply (7+sqrt 2) (8-3 sqrt 2)?

OpenStudy (anonymous):

Yes, I can help.

OpenStudy (anonymous):

Do you have any idea of what to do? Or where you might possible begin?

OpenStudy (anonymous):

My first reaction is to do foil method.

OpenStudy (anonymous):

And you would be correct.

OpenStudy (anonymous):

So i start with 7 *8 then \[7 * 3\sqrt{2}\]?

OpenStudy (cwrw238):

second part is 7 * -3sqrt2

OpenStudy (anonymous):

When im done with foil what do i do?

OpenStudy (anonymous):

*sigh* (7+sqrt 2) (8-3 sqrt 2) \[(7+\sqrt{2}) (8-\sqrt{2})\] \[56+8\sqrt{2}-21\sqrt{2} -3(2)\] \[56-13\sqrt{2}-6\] \[50-13\sqrt{2}\]

OpenStudy (anonymous):

So after foil just combine every thing?